OpenStudy (anonymous):

Jenny traveled to the lake and back to the trip there she traveled 80km/h and on the return she went 60km/h. How long did the trip there take if the return trip took 4 hours?

OpenStudy (anonymous):

neeed help

OpenStudy (anonymous):

s*t=d 60*4=240km time taken =240/80 =3hrs

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

but how you know 240/ 80?